A spring 20cm long is stretched to 25cm by a load of 50N. What will be it's length when stretched by 100N assuming that the elastic limit is not reached

1
Expert's answer
2022-03-04T08:30:52-0500

The Hooke's law says

"F=k\Delta l"

Hence

"l_2=\frac{F_2}{F_1}(l_1-l_0)+l_0\\
=100/50*(25-20)+20=30\:\rm cm"